OpenStudy (anonymous):

Verify the identity. sin 4x = 2 sin 2x cos 2x

OpenStudy (anonymous):

sin(4x) = sin( 2x + 2x) Recall that sin( A +B) = sin(A)cos(B) + cos(A)sin(B) so it is = sin2xcos2x + sin2xcos2x

OpenStudy (anonymous):

can you figure out the rest?

OpenStudy (anonymous):

What else do i do?

OpenStudy (anonymous):

what else is there to do other than add it ?

OpenStudy (anonymous):

I dont know!! haha

OpenStudy (anonymous):

haha well did I verify it ..now your done!

OpenStudy (anonymous):

thank you so extremely much. :)

OpenStudy (anonymous):

yw =))

OpenStudy (anonymous):

U can use the identity sin2y=2sinycosy!! and take sin4x=sin2(2x) where take 2x as y and prove it!!